OpenStudy (anonymous):

solve -9h-6+12h+40=22

OpenStudy (anonymous):

3h= -12 h= -4

OpenStudy (anonymous):

-9h-6+12h+40=22 -9h + 12h = 22-6-40 3h=-12 h=-12/3 = -4